OpenStudy (anonymous):

how do I convert r^2 sin(2 theta) = 4 to Cartesian form

OpenStudy (mr.math):

Hint: \(\sin(2\theta)=2\sin(\theta)\cos(\theta).\)

OpenStudy (anonymous):

you mean find r (the vlue of r given the expression) ?

OpenStudy (mr.math):

Is my hint good enough?

OpenStudy (anonymous):

@mr math, yeah, i think . .

OpenStudy (anonymous):